Maria is 36 years old. Maria was twice as old as Adrian was when Maria was as old as Adrian is now. How old is Adrian?

Answer by Theo(13342)

m = maria's age now.
a = adrian's age now.

maria is 36 years old now, so the formula for that is m = 36.

maria was twice as old as adrian was when maria was as old as adrian is now.
let x = the number of years ago when this occurred.

if you let x equal the number of years ago when maria was as old as adrian is now, then the formula for that is m - x = a

the statement says that maria was twice as old as adrian was when maria was as old as adrian is now.
the formula for that is m - x = 2 * (a - x)
since (m-x) = a, then replace a in this formula to get:
m - x = 2 * (m - x - x)
combine like terms to get:
m - x = 2 * (m - 2x)
simplify to get m - x = 2m - 4x
subtract m from both sides of the equation and add 4x to both sides of the equation to get 4x - x = 2m - m
combine like terms to get 3x = m

since m = 36, then you get:
3x = 36
solve for x to get:
x = 12

since m - 14 = a, then we also know that a is equal to 24.

your original formulas were:
m = 36
m - x = a
m - x = 2 * (a - x)

when m = 36 and x = 12, the formulas become:
m = 36 stays the same
m - x = a becomes 36 - 12 = 24 which becomes 24 = 24 which is true.
m - x = 2 * (a - x) becomes 36 - 12 = 2 * (24 - 12) which becomes 24 = 24 which is also true.

the values of m = 36 and x = 12 and a = 24 are good.

your solution is that adrian is 24 years old.